# DiffPF: Differentiable Particle Filtering with Generative Sampling via Conditional Diffusion Models

### v1

@@ V CONCLUSIONS
In this work, we proposed DiffPF, a novel differentiable particle filtering framework that integrates diffusion models into the recursive Bayesian estimation pipeline. Our method replaces traditional update mechanisms—typically reliant on hand-designed proposals and importance weighting—with a learned denoising process that enables flexible, fully differentiable, and high-quality sampling from complex posteriors. By modeling the update step with a conditional diffusion model, DiffPF is capable of capturing multimodal distributions, mitigating particle degeneracy, and eliminating the need for resampling. This design leads to more expressive and robust belief representations compared to existing differentiable particle filters. We validate the effectiveness of DiffPF across a variety of simulated and real-world tasks. Experimental results consistently show superior state estimation accuracy over strong baselines, demonstrating the generality and robustness of our approach.
Future work will focus on optimizing inference speed to better suit real-time applications and extending the framework to leverage richer sensory inputs and global correction techniques such as loop closure.

### v2

@@ V CONCLUSIONS
In this work, we introduced DiffPF, a novel differentiable particle filtering framework that integrates diffusion models into the recursive Bayesian estimation pipeline. DiffPF replaces hand-designed proposals and importance weighting with a learned denoising process, enabling flexible and fully differentiable sampling from complex posteriors. This conditional diffusion formulation effectively captures multimodal distributions, mitigates particle degeneracy, and removes the need for resampling, yielding more expressive and robust belief representations. Experiments on both simulated and real-world tasks demonstrate consistent performance gains over strong baselines, highlighting the generality and robustness of our approach. Future work may focus on the potential risk of sample collapse and mode dropping by exploring weighted variants of DiffPF, where importance sampling could further improve robustness.